Resolving to read more each week, Drew starts a new reading program and promises to read 10 minutes the first week, 20 minutes the second, 30 minutes the third, and so on.
How much time will Drew spend reading during the 12th week?
After one year, how much time did Drew spend reading?
1
Expert's answer
2020-04-07T16:05:24-0400

During n-th week Drew spend 10*n minutes reading.


During 12th week = 10 * 12 = 120 minutes.


Year has 52 weeks. After one year Drew spend reading


"\quad \displaystyle\sum_{n=1}^{52} 10*n = 10 * \dfrac{(1+52)*52}{2} = 13,780 \text{ minutes} \approx 230 \text{ hours}"